OpenStudy (anonymous):

Find equations of the lines through the given point parallel to the given line and perpendicular to the given line. Point (7/8, 3/4) Line 5x + 3y = 0

OpenStudy (anonymous):

This is just really confusing me right now :/

OpenStudy (anonymous):

slope the line parallel to 5x+ 3y = 0 is given by m = -5/3 then using the formula y = mx + c find the eqn , use the above slope and substitute x,y = (7/8,3/4) for perpendicular line slope = -1/slope = 3/5 use the eqn y = mx + c substitute the m value and x,y values

OpenStudy (anonymous):

alrighty, let me try this

OpenStudy (anonymous):

did u get the answer??

OpenStudy (anonymous):

still working on it haha sorry im slow

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i think i may still be doing it wrong...

OpenStudy (anonymous):

is it okay if i convert to decimals?

OpenStudy (anonymous):

ok if your comfortable with using decimals you can use it... did find the slope ??

OpenStudy (anonymous):

is it y= -2.2083?

OpenStudy (anonymous):

you shouldnt find y first take the case of parallel line tell me the slope ??

OpenStudy (anonymous):

-5/3

OpenStudy (anonymous):

ok substitute the slope and x,y values in the equation y = mx + c to find c

OpenStudy (anonymous):

3/4=-5/3(7/8)+c

OpenStudy (anonymous):

find c

OpenStudy (anonymous):

c= 2.654761905

OpenStudy (anonymous):

ok. substitute the slope and c value in the eqn y = mx + c to get the eqn of line parallel to 5x + 3y = 0

OpenStudy (anonymous):

y=-5/3x+2.65471905

OpenStudy (anonymous):

2.654761905*

OpenStudy (anonymous):

ok. your right now solve for perpendicular line

OpenStudy (anonymous):

m=3/5

OpenStudy (anonymous):

ok go ahead . follow same procedure and find the solution.

OpenStudy (anonymous):

y=3/5x+.225

OpenStudy (anonymous):

i think its correct... you followed same procedure right.....

OpenStudy (anonymous):

yes i did haha

OpenStudy (anonymous):

:)